Consider the following short run production function: Q=6L2-0.4L3 

1. a) Find the value of L that maximizes output

b) Find the value of L that maximizes marginal product

c) Find the value of L that maximizes average product


1
Expert's answer
2021-02-10T21:38:48-0500

(a) Let's first find the marginal product:


"MP=\dfrac{d}{dL}(Q)=\dfrac{d}{dL}(6L^2-0.4L^3)=12L-1.2L^2."

Then. we can find the value of L that maximizes output:


"MP=0,""12L-1.2L^2=0,""L=10."

b) Let's take the derivative from MP with respect to L:


"\dfrac{d}{dL}(MP)=\dfrac{d}{dL}(12L-1.2L^2)=12-2.4L."

Then, we can find the value of L that maximizes marginal product:


"12-2.4L=0,""L=5."

c) Let's find the AP:


"AP=\dfrac{Q}{L}=\dfrac{6L^2-0.4L^3}{L}=6L-0.4L^2."

Let's take the derivative from AP with respect to L:


"\dfrac{d}{dL}(AP)=\dfrac{d}{dL}(6L-0.4L^2)=6-0.8L."

Then, we can find the value of L that maximizes average product:


"6-0.8L=0,""L=7.5"